# TabICL: A Tabular Foundation Model for In-Context Learning on Large Data

### v1

@@ 4.1 Improved Pretraining Synthetic Datasets
As tree-based models excel on tabular data, we introduce tree-based SCMs to leverage their ability to model complex interactions and hierarchical dependencies between variables. We define $f$ using an XGBoost regression model, as it is widely favored by practitioners and supports multi-output regression. At each layer of the graph, an XGBoost model is trained on fake targets drawn from Gaussian noise, taking the values of the parent variables as input. The obtained predictions then become the values of the child variables. To balance data generation, we combine SCMs (70%) with tree-based SCMs (30%). Appendix B gives details and examples of generated data.
@@ 4.2 Curriculum Learning for Large-scale Pretraining
Similar to pretraining LLMs on shorter sentences before moving to longer ones, we gradually increase the size of synthetic datasets (i.e., the number of samples) while adjusting the micro batch size $N_{\mathcal{B}}$ used for gradient accumulation to accommodate memory constraints, as follows:
@@
    $N_{\mathcal{B}}=4$ with a fixed size of 1,024 for 100K steps;
@@
Each step consists of 512 datasets with the number of features ($\leq 100$) and classes ($\leq 10$) randomly sampled. FlashAttention and automatic mixed precision are applied globally. The pretraining took 2 weeks on three A100 GPUs with 40GB memory using PyTorch (10, 3, and 1 days for stage 1, 2, and 3, respectively). Section D.1 gives more pretraining details.
@@ 4.3 Hierarchical Class-extension Strategy
We tackle many-class classification problems ($>10$ classes) through hierarchical classification (Silla & Freitas, 2011). Specifically, we recursively and evenly partition classes into subgroups of up to 10 classes, forming a multi-level classification tree. A classification problem with $k$ classes requires a hierarchy with depth $r=\left\lceil\log_{10}k\right\rceil$. Each node in the tree corresponds to a sub-task that predicts probabilities for its subgroup. During inference, the final probability for a given class is obtained by multiplying the probabilities across all relevant nodes from root to leaf.
As noted earlier, labels are used exclusively in the final ICL block. Consequently, the hierarchical tree is constructed during dataset-wise ICL, with all sub-tasks sharing the learned row embeddings $H$ and the same $\text{TF}_{\text{icl}}$ for ICL predictions. This kind of sharing greatly enhances the efficiency of TabICL in hierarchical classification scenarios.
@@ 4.4 Memory-efficient Inference
Using FlashAttention, which offers linear memory complexity with respect to sequence length, we observed that the peak activation memory can be effectively modeled by a polynomial regression during inference: $\alpha_{1}\times\text{batch\_size}+\alpha_{2}\times\text{seq\_len}+\alpha_{3}\times\text{batch\_size}\times\text{seq\_len}+\alpha_{4}$. This enables us to dynamically adjust the batch size based on sequence length and available GPU memory. The batch dimension serves different roles upon the context: it represents the number of columns for column-wise embedding, the sample size for row-wise interaction, and the number of datasets for dataset-wise ICL. Inspired by Rajbhandari et al. (2021), intermediate activations can be offloaded to CPU and disk as needed to further reduce memory consumption. These optimizations enable TabICL to handle datasets with 100K samples and 500 features using only 5 GB of GPU memory and 32 GB of RAM. This suffices for most real-world applications. See Section D.2 for details.
@@ 5.1 Benchmark
We evaluate TabICL on the benchmark by Ye et al. (2025), referred to as “TALENT” here. It comprises 200 classification datasets (120 binary and 80 multiclass datasets) and comes with results for over 30 baselines. We will first analyze the 188 datasets with at most 10 classes, since these can be handled natively by TabICL and TabPFNv2.
Datasets are split into 64% training, 16% validation, and 20% test data. While most models rely on the validation set for early stopping and hyperparameter tuning, TabICL and TabPFNv2 do not. Nonetheless, we opted to train TabICL and TabPFNv2 using only the training data, which places them at a disadvantage. On the flip side, both TabICL and TabPFNv2 leverage ensembles, unlike the other deep learning models. A fair comparison would require training the other models as an ensemble as well. However, this is computationally expensive, was not part of the original benchmark, and is not implemented in these experiments. Both TabICL and TabPFNv2 average over 32 predictions obtained by randomly shuffling columns and classes and using different pre-processors. TabICL applies $z$-normalization with or without power transformation to all input features. See Hollmann et al. (2025) for details on TabPFNv2. To avoid out-of-memory issues, we subsample training sets to 30K samples for TabPFNv2, while TabICL could predict on all datasets without subsampling. In addition, we disable automatic mixed precision across all methods during inference to ensure reproducibility and consistent time measurements.
@@ 5.2 Results
Figure 5: Accuracy and training/inference times on the TALENT benchmark up to 10 classes. For TabICL and TabPFNv2, the time reported corresponds to training+inference on an A100 GPU. For the other models, it includes both training and hyperparameter tuning. Since Ye et al. (2025) track only the training time using the best hyperparameters found after 100 tuning steps, we approximate the total training time by multiplying this value by 100. For TabPFN (v1), we do not report the time since the original time measurement does not include the inference time.
@@
Figure 5 shows accuracies relative to the tuned MLP for all models. TabICL obtains the best median relative accuracy across all datasets while being much faster than traditional state-of-the-art models: the geometric mean training+inference time per 1K samples for TabICL is 1.1 seconds, while tuning CatBoost on a CPU takes around 3 minutes, and RealMLP and ModernNCA on a GPU take around 7 minutes. Looking at the ranks of each method based on the obtained accuracies, the critical difference diagram in Figure E.1 shows that TabICL and TabPFN2 outperform competitors by a wide margin, while the difference between the two of them is not statistically significant.
@@
Figure 6 shows that TabICL is 1.5$\times$ faster than TabPFNv2 on small datasets and 3-10$\times$ faster on large datasets. This is facilitated by the hybrid architecture of TabICL, using fewer of the expensive row-wise and column-wise attention layers with smaller embedding dimension before ICL on tokenized rows. On a dataset with 10,000 samples and 100 features, TabICL takes roughly 20 seconds versus 1 minute 40 seconds for TabPFNv2, while for a dataset of 1000 samples and 10 features, TabICL takes 1 second compared to 2 seconds for TabPFNv2. In Figure A.1, we fit simple scaling laws to predict the runtime of TabICL and TabPFN2 based on the number of samples and features. These scaling laws show that the average speedup of TabICL remains around five-fold for large datasets.
@@
While TabPFNv2 achieves excellent performance on datasets up to 10K samples, it has only been pre-trained with up to 2048 training samples and can fail on datasets above 30K samples due to its memory usage. Figure 7 shows that unlike TabPFNv2, the performance of TabICL remains strong for larger datasets. While ICL is often used in few-shot settings, this demonstrates that foundation models can compete with the best deep learning or tree-based models, even in large-sample regimes where the latter have the most potential. Additional results in Appendix A demonstrate that TabICL also performs well for many classes, features, and high ratios of categorical features.
@@
In many practical situations, the quality of the estimated probabilities is crucial for decision-making. Therefore, we also report the log loss (a.k.a. cross-entropy loss), which is a proper scoring rule and therefore rewards accurate prediction of probabilities (Gneiting & Raftery, 2007). Since TabICL does not leverage hyperparameter tuning, its predictions are not specifically optimized towards accuracy. The critical difference diagrams in Appendix E show that TabICL and TabPFN2 significantly outperform accuracy-tuned competitors on the log loss, though the difference between the two is not significant, indicating that both produce more reliable probability estimates than the other models.
@@ 6 Conclusion
We introduce TabICL, a novel tabular foundation model that extends the scalability of existing tabular foundation models by an order of magnitude. Evaluated on datasets with up to 100K training samples, it delivers excellent performance without hyperparameter tuning, making it nearly two orders of magnitude faster than other tabular methods requiring hyper-parameter tuning. TabICL achieves this through a hybrid architecture and memory-saving optimizations. Compared to the newly released and leading tabular foundation model — TabPFNv2, our TabICL achieves comparable performance while being more scalable and faster.
@@
Like other foundation models, TabICL suffers from slow inference speed, although TabPFNv2 has demonstrated that this issue can be alleviated to some extent through caching. Currently, TabICL is limited to classification problems, but as Hollmann et al. (2025) showed, regression problems can be treated with similar methodology. Our evaluation inherits the strengths and weaknesses of the TALENT benchmark. In particular, TALENT as most other benchmarks trains single models tuned using the holdout method, while ensembling models evaluated using cross-validation can improve performance. Bagging increases computational cost, whereas TabICL can simply be trained on the full data without needing a validation set.
@@
In-context learning for tabular data was originally introduced as a speedup for small tables (Hollmann et al., 2022). The expressive power of a forward pass in a transformer may seem limited, and one might wonder if in-context learning loses its edge, given sufficient data. We find that even with large data, pre-training can create implicit priors that give in-context transformers a competitive advantage.
@@ Acknowledgements
We thank Léo Grinsztajn for coming up with the idea to add a tree-based part to the prior. We also thank Lennart Purucker and Samuel Müller for interesting discussions.
Appendix A Further Experiments

### v2

@@ 3.4 Dataset-wise In-Context Learning
3.5 Computational Complexity Analysis
To characterize the computational efficiency and scalability of TabICL, we analyze the time complexity of its three components for a table with $n$ rows and $m$ features:
  1. 1.
    $\text{TF}_{\text{col}}$ processes each column independently using ISAB, resulting in complexity $\mathcal{O}(nkm)$, where $k$ is the fixed number of inducing vectors.
  2. 2.
    $\text{TF}_{\text{row}}$ applies self-attention over features and 4 [CLS] tokens per row, yielding complexity $\mathcal{O}(m^{2}n)$. $\text{TF}_{\text{row}}$ is efficient since $m\ll n$ in most tabular settings.
  3. 3.
    $\text{TF}_{\text{icl}}$ operates on the compressed row embeddings via self-attention with complexity $\mathcal{O}(n^{2})$. While quadratic in $n$, its cost is incurred once since the feature dimension is collapsed, ensuring practical efficiency.
The overall complexity of TabICL is $\mathcal{O}(m^{2}n+n^{2})$. In contrast, TabPFNv2 alternates column-wise and row-wise attentions without collapsing dimensions, resulting in complexity $\mathcal{O}(m^{2}n+n^{2}m)$. As a result, TabPFNv2 can become computationally expensive for large $n$ and moderate $m$.
@@ 4.1 Improved Pretraining Synthetic Datasets
Inspired by Grinsztajn (2024), we introduce tree-based SCMs to leverage their ability to model complex interactions and hierarchical dependencies between variables. We define $f$ using an XGBoost regression model, as it is widely favored by practitioners and supports multi-output regression. At each layer of the graph, an XGBoost model is trained on fake targets drawn from Gaussian noise, taking the values of the parent variables as input. The obtained predictions then become the values of the child variables. To balance data generation, we combine SCMs (70%) with tree-based SCMs (30%). Appendix C gives details and examples of generated data.
@@ 4.2 Curriculum Learning for Large-scale Pretraining
Similar to pretraining LLMs on shorter sentences before moving to longer ones, we gradually increase the size of synthetic datasets (i.e., the number of samples) while adjusting the micro batch size $N_{\mathcal{B}}$ used for gradient accumulation to accommodate memory constraints. We employed a three-stage procedure:
@@
    $N_{\mathcal{B}}=4$ with a fixed size of 1,024 for 160K steps;
@@
Each step consists of 512 datasets with the number of features ($\leq 100$) and classes ($\leq 10$) randomly sampled. FlashAttention and automatic mixed precision are applied globally. The pretraining took 20 days on three A100 GPUs with 40GB memory using PyTorch (16, 3, and 1 days for stage 1, 2, and 3, respectively). Section E.1 gives more pretraining details.
@@ 4.3 Hierarchical Class-extension Strategy
We tackle many-class classification problems ($>10$ classes) through hierarchical classification (Silla & Freitas, 2011). Specifically, we recursively and evenly partition classes into subgroups of up to 10 classes, forming a multi-level classification tree. A classification problem with $k$ classes requires a hierarchy with depth $r=\left\lceil\log_{10}k\right\rceil$. Each internal node in the tree predicts the probability of its child groups, and each leaf predicts the probability of the classes it contains. During inference, the final probability for a given class is obtained by multiplying the probabilities across all nodes situated on the unique root-to-leaf path leading to this class. See Section E.3 for more details.
As noted earlier, labels are used exclusively in the final ICL block. Consequently, the hierarchical tree is constructed during dataset-wise ICL, with all sub-tasks sharing the same learned row embeddings $H$ and the same $\text{TF}_{\text{icl}}$ for ICL predictions. This kind of sharing greatly enhances the efficiency of TabICL in hierarchical classification scenarios.
@@ 4.4 Memory-efficient Inference
Using FlashAttention, which offers linear memory complexity with respect to sequence length, we observed that the peak activation memory can be effectively modeled by a polynomial regression during inference: $\alpha_{1}\times\text{batch\_size}+\alpha_{2}\times\text{seq\_len}+\alpha_{3}\times\text{batch\_size}\times\text{seq\_len}+\alpha_{4}$. This enabled us to dynamically adjust the batch size based on sequence length and available GPU memory. The batch dimension serves different roles depending on the context: it represents the number of columns for column-wise embedding, the sample size for row-wise interaction, and the number of datasets for dataset-wise ICL. Inspired by Rajbhandari et al. (2021), intermediate activations can be offloaded to CPU and disk as needed to further reduce memory consumption. These optimizations enable TabICL to handle datasets with 100K samples and 500 features using only 5 GB of GPU memory and 32 GB of RAM. This suffices for many real-world applications. See Section E.2 for details.
@@ 5.1 Benchmark
We evaluate TabICL on the TALENT benchmark introduced by Ye et al. (2025). It comprises 200 classification datasets (120 binary and 80 multiclass datasets) and comes with results for over 30 baseline methods. To ensure a fair comparison, we exclude 15 datasets that were used by TabPFNv2 for hyperparameter tuning and model selection (see Appendix F for more details). In what follows, we mainly focus on the 171 datasets with at most 10 classes, since these can be handled natively by TabICL and TabPFNv2.
Datasets are split into 64% training, 16% validation, and 20% test data. While most models rely on the validation set for early stopping and hyperparameter tuning, TabICL and TabPFNv2 do not. Nonetheless, we opted to train TabICL and TabPFNv2 using the training data only, which places them at a disadvantage. On the flip side, both TabICL and TabPFNv2 leverage ensembles, unlike the other deep learning models. In Appendix H, we show that increasing ensemble size improves performance for both TabPFNv2 and TabICL, with TabPFNv2 benefiting more from ensembling. A fair comparison would require training the other models as an ensemble as well. However, this is computationally expensive, was not part of the original benchmark, and is not implemented in these experiments.
Both TabICL and TabPFNv2 average over 32 predictions obtained by randomly shuffling columns and classes and using different pre-processors. TabICL applies $z$-normalization with or without power transformation to all input features. See Hollmann et al. (2025) for details on TabPFNv2. To avoid out-of-memory issues, we subsample training sets to 30K samples for TabPFNv2, while TabICL could predict on all datasets without subsampling. In addition, we disable automatic mixed precision across all methods during inference to ensure reproducibility and consistent time measurements.
@@ 5.2 Results
Figure 5 shows accuracies relative to the tuned MLP for all models. TabICL obtains the best median relative accuracy across all datasets while being much faster than traditional state-of-the-art models: the geometric mean training+inference time per 1K samples for TabICL is 1.1 seconds, while tuning CatBoost on a CPU takes around 3 minutes, and RealMLP and ModernNCA on a GPU take around 7 minutes. Looking at the ranks of each method based on the obtained accuracies, the critical difference diagram in Figure I.1 shows that TabICL and TabPFNv2 outperform competitors by a wide margin, while the difference between the two of them is not statistically significant.
Figure 5: Accuracy and training/inference times on the TALENT benchmark up to 10 classes. The red bar indicates the median relative accuracy improvements over MLP across all datasets for each method. For TabICL and TabPFNv2, the time reported corresponds to training+inference on an A100 GPU. For the other models, it includes both training and hyperparameter tuning. Since Ye et al. (2025) track only the training time using the best hyperparameters found after 100 tuning steps, we approximate the total training time by multiplying this value by 100. For TabPFN (v1), LoCalPFN, and TuneTables, we do not report the time since the original time measurement does not include the inference time.
@@
Figure 6 shows that TabICL is 1.5$\times$ faster than TabPFNv2 on small datasets and 3-10$\times$ faster on large datasets. This is facilitated by the hybrid architecture of TabICL, using fewer of the expensive row-wise and column-wise attention layers with smaller embedding dimension before ICL on tokenized rows. On a dataset with 10,000 samples and 100 features, TabICL takes roughly 20 seconds versus 1 minute 40 seconds for TabPFNv2, while for a dataset of 1000 samples and 10 features, TabICL takes 1 second compared to 2 seconds for TabPFNv2. The empirical speedup of TabICL over TabPFNv2 also corroborates the complexity analysis in Section 3.5. In Figure B.1, we fit simple scaling laws to predict the runtime of TabICL and TabPFNv2 based on the number of samples and features. These scaling laws show that the average speedup of TabICL remains around five-fold for large datasets.
@@
While TabPFNv2 achieves excellent performance on datasets up to 10K samples, it has only been pre-trained with up to 2048 training samples and can fail on datasets above 30K samples due to its memory usage. Figure 7 shows that unlike TabPFNv2, the performance of TabICL remains strong for larger datasets. While ICL is often used in few-shot settings, this demonstrates that foundation models can compete with the best deep learning or tree-based models, even in large-sample regimes where the latter have the most potential. Additional results in Appendix B demonstrate that TabICL also performs well for many classes, features, and high ratios of categorical features. To enhance the performance of TabPFNv2 on large datasets, Hollmann et al. (2025) proposed a random forest extension: during training, the data is partitioned using decision trees, and TabPFNv2 is fitted at each leaf to enable localized predictions at inference time. We found that this extension can significantly improve the performance of both TabPFNv2 and TabICL on large datasets (refer to Appendix G for more details).
@@
In many practical situations, the quality of the estimated probabilities is crucial for decision-making. Therefore, we also report the log loss (a.k.a. cross-entropy loss), which is a proper scoring rule and therefore rewards accurate prediction of probabilities (Gneiting & Raftery, 2007). Since TabICL does not leverage hyperparameter tuning, its predictions are not specifically optimized towards accuracy. The critical difference diagrams in Appendix I show that TabICL and TabPFNv2 significantly outperform accuracy-tuned competitors on the log loss, though the difference between the two is not significant, indicating that both produce more reliable probability estimates than the other models.
@@
5.3 Ablation Studies
Tree-based SCMs prove beneficial.
Incorporating the inductive biases of tree-based models during pretraining improve the performance of TabICL, as shown in Figure 9.
Figure 9: Relative improvements of TabICL pretrained on prior datasets with versus without tree-based SCMs across 200 datasets. Green triangles denote the mean relative improvements. To mitigate computational overhead, we pretrained TabICL for 20K steps following the stage 1 setup of curriculum learning.
Curriculum learning effectively improves the performance of TabICL on large datasets.
Across the three stages of curriculum learning, the average rank of TabICL improves from 11.4 (ninth place), to 7.46 (second), and finally to 6.95 (first), as shown in Figure E.2. These results show the effectiveness of curriculum learning. However, we also observe a slight performance drop on some small datasets during curriculum learning, as shown in Figure 10.
Figure 10: Relative accuracy improvements over Stage 1 across 200 classification datasets. Green triangles denote the mean relative improvements.
@@ 6 Conclusion
We introduce TabICL, a novel tabular foundation model that extends the scalability of existing tabular foundation models by an order of magnitude. Evaluated on datasets with up to 100K training samples, it delivers excellent performance without hyperparameter tuning, making it nearly two orders of magnitude faster than other tabular methods requiring hyper-parameter tuning. TabICL achieves this through a hybrid architecture and memory-saving optimizations. Compared to the newly released and leading tabular foundation model TabPFNv2, TabICL achieves comparable performance while being more scalable and faster.
@@
(1) Like other foundation models, TabICL suffers from slow inference speed, though TabPFNv2 has shown caching can alleviate this. Currently, TabICL is limited to classification, but regression can be treated with similar methodology, as Hollmann et al. (2025) showed. (2) A defining characteristic of tabular data is that columns do not have a natural ordering. Ideally, tabular methods should be invariant to column permutations. However, TabICL violates this invariance due to the use of RoPE. (3) Our evaluation inherits the strengths and weaknesses of the TALENT benchmark. Notably, TALENT trains single models tuned using the holdout method, while ensembling models evaluated using cross-validation can improve performance. Bagging increases computational cost, whereas TabICL can simply be trained on the full data without needing a validation set. (4) TALENT employs mean imputation for handling missing values. However, it might be better to allow TabPFNv2 to handle missing values internally despite the low proportion of missing values in TALENT. (5) We did not provide categorical information to TabPFNv2. Though TabPFNv2 infers categorical features internally, providing this information explicitly could enhance its performance.
@@
In-context learning for tabular data was originally introduced as a speedup for small tables (Hollmann et al., 2022). As the expressive power of a forward pass in a transformer may seem limited, one might wonder if in-context learning loses its edge given sufficient data. We find that even with large data, pre-training induces implicit priors that give in-context transformers a competitive advantage.
@@ Acknowledgements
We thank Lennart Purucker and Samuel Müller for interesting discussions. We are also grateful to Si-Yang Liu and Han-Jia Ye, the authors of the TALENT benchmark, for providing information and clarifying our questions. Finally, we thank the anonymous ICML reviewers for their valuable feedback and suggestions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Systematic Comparison between TabICL and TabPFNv2
This appendix provides a systematic comparison between TabICL and TabPFNv2, two tabular foundation models that leverage In-Context Learning (ICL) to make predictions. Both models are pretrained on synthetic datasets and aim to deliver strong performance on unseen real-world tabular datasets.
A.1 Architecture
Core mechanism.
(1) TabICL employs a two-stage architecture, where the first stage uses a column-then-row attention mechanism to build fixed-dimensional embeddings and effectively collapses the column dimension to reduce computational complexity for the subsequent ICL stage. (2) TabPFNv2 utilizes a two-way attention mechanism that alternates between column-wise and row-wise attentions. These attentions operate on the input table across original dimensions without any intermediate collapse.
Representation collapse.
When features have similar distributions, both TabICL and TabPFNv2 may suffer from the representation collapse issue. (1) To alleviate this problem, TabICL incorporates Rotary Positional Embedding (RoPE) into the transformer used in row-wise interaction. RoPE helps break the symmetry between identically distributed features by encoding relative positional information. (2) TabPFNv2 addresses this by introducing random feature identifier vectors for each feature. It also encodes groups of features collectively rather than individually.
Label fusion.
TabICL and TabPFNv2 differ significantly in when they incorporate label information. (1) TabICL employs a late fusion strategy. TabICL processes input features independently of their labels during row-wise embedding and column-wise interaction. Labels are utilized exclusively in the final ICL stage. (2) TabPFNv2 employs an early fusion strategy for label information. TabPFNv2 directly concatenates the input features with the labels of the training samples, processing them together from the very beginning and throughout its layers.
A.2 Pretraining
Prior generation.
Both TabICL and TabPFNv2 are pretrained exclusively on synthetic prior datasets. (1) TabICL builds on the previous prior generation of TabPFN (v1), which relies on Structural Causal Models (SCMs). TabICL introduces two more extensions: broader nonlinear modeling capabilities of SCMs and the integration of tree-based SCMs to incorporate the inductive biases of tree-based models. (2) TabPFNv2 constructs diverse Directed Acyclic Graphs (DAGs) through a sophisticated procedure (i.e., growing networks with redirection sampling method) and incorporates various computational mappings at graph edges, including small neural networks, categorical feature discretization, and decision trees. However, the code for its prior generation is not open-source.
Dataset size.
(1) TabICL was pretrained using curriculum learning that progressively scales the number of samples from 1,024 to 60K. The number of features was sampled uniformly up to 100, and the number of classes was sampled uniformly up to 10. (2) TabPFNv2 was pretrained on synthetic datasets containing up to 2,048 training samples (uniformly sampled). The number of features was drawn from a beta distribution and then linearly scaled to the range from 1 to 160, while the number of classes was also uniformly sampled up to 10.
Pretraining scale.
(1) TabICL was pretrained over approximately 82 million synthetic datasets. (2) TabPFNv2 was trained on a larger corpus, comprising approximately 130 million synthetic datasets.
A.3 Scalability and Efficiency
Dataset size handling.
(1) At inference, TabICL is capable of handling tables with 100K samples and 500 features within 5GB of GPU memory, and can deal with any number of classes using hierarchical class-extension strategy. (2) TabPFNv2 excels on tables with up to 10K samples and 500 features, and can only deal with up to 10 classes.
Computational complexity.
Given a table with $n$ rows and $m$ columns, the time complexity of TabICL is $\mathcal{O}(m^{2}n+n^{2})$, while the time complexity of TabPFNv2 is $\mathcal{O}(m^{2}n+n^{2}m)$. Our extensive experiments demonstrate that TabICL is systematically faster than TabPFNv2.
Appendix B Further Experiments